OpenStudy (anonymous):

lim x=1 2-2x(square)/ x-1= ?

OpenStudy (anonymous):

\[\lim_{x\rightarrow 1} \frac{2-2x^2}{x-1}\]\[=\lim_{x\rightarrow 1} \frac{2(1-x^2)}{x-1}\]\[= \lim_{x\rightarrow 1} \frac{2(1-x)(1+x)}{x-1}\]\[=\lim_{x\rightarrow 1} \frac{-2(x-1)(x+1)}{x-1}\] And then cancel the x-1 factors.

OpenStudy (anonymous):

and plug in 1 for x

OpenStudy (anonymous):

can we just plug 1 in directly

OpenStudy (anonymous):

no, because you will get 0/0, an indeterminant solution

OpenStudy (anonymous):

thanks to both of you